i worked this problem: (2x + 1)/3 - (3-x)/4 = (x)/6
and i got 1.67 = x which is wrong
this is how i did it:
(2x+1)/3 - (3-x)/4 = (x)/6
multiply by the LCD,12. to get rid of the fractions
24x+12/3 - 36-12x/4 = 12x/6
8x + 4 - 9 - 3x = 2x
simplify
5x - 5 = 2x
subtract 2x to get the variables on the same side
5x - 5 - 5x = 2x - 5x
-5 = -3x
multiply by 1/-3
-5(1/-3) = -3x(1/-3)
-5/-3 = -3x/-3
simplify
1.67 = x
can you tell me what am i doing wrong??? thanks!











Answer by jrfrunner(365)

following your step
multiply by the LCD,12. to get rid of the fractions
--
you did
24x+12/3 - 36-12x/4 = 12x/6
you are not clear with how you are multiplying both sides
(24x+12)/3 -(36-12x)/4=12x/6
--
which yiels
8x+4-9+3x=2x you got 8x+4-9-3x=2x you have missed distributing the negative to -12x
===
by the way, it is much easier to do the following, after you multiply the LCD
12(2x+1)/3-12(3-x)/4=12x/6
after this step, simply the ratios
4(2x+1)-3(3-x)=2x